The ratio of the magnetic field inside a solenoid at an axial point well inside and at an axial end point is

Ans 2


1
Expert's answer
2020-11-17T11:19:04-0500

if magnetic permeability ="\mu" , current flowing =I ,Number of turns per unit length =n

then magnetic field inside solenoid = "\mu \times n\times I" .............(1)

Magnetic field near the Axial End point = "\frac{\mu \times n \times I}{2}" ............(2)

Hence field ratio = "\frac{equation 1}{equation 2} = 2"

Answer =2